A regulation soccer field is 50 yd wide and 100 yd long. Find the length of the diagonal of the field.
Answer by checkley77(12844)

50^2+100^2=c^2
2,500+10,000=c^2
12,500=c^2
c=sqrt12,500
c=111.8 yds. ans, for the length of the diagonal.